OpenStudy (anonymous):

I was absent from school the other day and I need someone to help explain how to do some math problems.... (4/2x)=(7/3) please, if you could, explain how you got your answer. Thanks!

OpenStudy (inkyvoyd):

Do you mean \(\huge \frac{4}{2x}=\frac{7}{3}\) or \(\huge \frac{4}{2}*x=\frac{7}{3}\)

OpenStudy (anonymous):

You need get x alone. What you do to one side, you must also do to the other. This means that you can divide both sides by 4/2 to find the value of x.

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

so you would do (7/3) / (4/2)?

OpenStudy (alexwee123):

you would cross multiply

OpenStudy (alexwee123):